OpenStudy (anonymous):

need steps please....x squared+x+5=0. Solve by completing the square

OpenStudy (dumbcow):

take middle coefficient, halve it, then square it x^2 +x +1/4 +5 -1/4 =0 (x+1/2)^2 = -19/4 x +1/2 = +-sqrt19i /2 x = (+-sqrt19i -1)/2

OpenStudy (anonymous):

Thank u....one more pls